# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

### v1

@@ Abstract
The appearance of a black hole shadow, in both astronomical observations and theoretical analysis, depends critically on the properties of the surrounding accretion. In this study, we analyse the emission and observational properties of a Schwarzschild black hole surrounded by a Hernquist dark matter (DM) halo by considering three simplified accretion models, which have not been previously explored in such detail. Based on our findings, direct emission contributes markedly to the whole observed intensity for a Schwarzschild-Hernquist black hole encircled by a geometrically thin accretion disk. The corresponding regions and measured intensities of the lensing ring, the photon ring, and direct emission are sensitive to the parameters of the Hernquist DM halo. Apart from disk accretion, we also study two spherical accretion models: static and infalling. With increasing Hernquist DM parameters, both models exhibit a significant decrease in the measured intensity, yet the photon spheres are enlarged by an increase of $\sim 2\%$ to $\sim 30\%$. Furthermore, Doppler de-boosting due to the infalling accretion results in darker shadow images compared to static cases. By precising visualisations of the shadows, our results suggest that the observational characteristics of black hole shadows can serve as a powerful probe for distinguishing between different DM distributions around supermassive black holes.
@@ I Introduction
The observation of black hole shadows has entered a new era, establishing them as a ground-breaking tool for researching powerful gravitational forces. The fundamental nature of gravity has been further illuminated by analysis of the critical behaviour of photons close to black holes. Recent advances in the Event Horizon Telescope’s (EHT) imaging of the supermassive black holes $\text{M87}^{*}$ Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023) and $\text{SgrA}^{*}$ Ghez et al. (2005); Boehle et al. (2016); Akiyama et al. (2022a, b, c, d, e, f, 2024a, 2024b) have shown shadow features that not only support general relativity (GR) Hawking (1972); Bardeen et al. (1973a); Wald (1994, 2010) predictions but also provide a novel window of probing black holes. The black hole shadow, also known as the photon capture zone, is the centre dark area shown in these pictures Perlick and Tsupko (2022). This is encircles a compact, asymmetrical and dazzling circular structure. This phenomenon, known as the photon ring, results from light beams from infinity being deflected by the strong gravitational field as they approach the black hole. In backward ray-tracing methods, light rays that approach the critical impact parameter asymptotically converge to the bound photon orbit. The so-called photon sphere, which projects the shadow of the black hole, is made up of this confined photon orbit. An angular radius equation for the shadow of Schwarzschild black holes has been established by Synge and Luminet using a crucial impact parameter of $b_{p}=3\sqrt{3}M$ Synge (1966); Luminet (1979). The first image of a black hole surrounded by a thin accretion disk was analytically produced by the authors in Ref.Luminet (1979), displaying primary and secondary images that emerge beyond the black hole’s shadow. Bardeen later studied the D-shaped shadow of Kerr black holes, showing that the shape of the shadow can be changed by the spin of the black hole Bardeen et al. (1973b); Chandrasekhar (1998). Additionally, a great deal of research has been done on black hole shadows in higher-dimensional spacetime and other modified gravity theories Amarilla et al. (2010); Amarilla and Eiroa (2012); Amir et al. (2018); Mizuno et al. (2018); Eiroa and Sendra (2018); Banerjee et al. (2020). In contrast to single shadows, the research has also suggested multiple shadows for wormholes and black holes Huang et al. (2016); Guerrero et al. (2021). While the finding of the EHT mainly relates to black holes in GR, it provides a great deal of opportunity to investigate other compact objects in theories other than GR.
@@
To investigate the main characteristics of black hole pictures and find possible indications of novel physics, it is helpful to set up a basic accretion model. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accretion settings the shadow has resilient properties where its shape and size are determined by spacetime instead of accretion details Gralla et al. (2019). Further investigations of optically thin and geometrically thin accretion models show that the size of the shadow is strongly related to the position of the accretion Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are always present in the bright areas outside the shadow. These characteristics make it possible to distinguish GR black holes from other compact objects or black holes in modified gravity theories through observation Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are extremely important for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@ II Light deflection and photon orbit
We utilises the Hernquist DM density distribution model to characterise a Schwarzschild black hole situated embedded a Hernquist DM haloDehnen (1993); Mo et al. (2010),
@@
The radius $r_{p}$ and the impact parameter $b_{p}$ of the photon sphere are represented in Eqs.(15)-(16), respectively. Tab.1 displays numerical results for the event horizon radius $r_{h}$, photon sphere radius $r_{p}$, and impact parameter $b_{p}$ by choosing various values for $\rho_{c}$ and $r_{s}$. It is evident that as $\rho_{c}$ and $r_{s}$ increase, the event horizon radius $r_{h}$, the photon sphere radius $r_{p}$, and the impact parameter $b_{p}$ all rise, exceeding their values in the Schwarzschild case.
@@ III.1 Direct emission, lensing ring and photon ring
According to Ref.Gralla et al. (2019), radiation with different impact parameters $b$ would cause the black hole to look differently to a distant observer when tracing rays backwards from an observer towards the black hole. We may use ray-tracing procedure to determine the optical appearance of a Schwarzschild-Hernquist black hole, by calculating the total deflection in azimuthal angle $\varphi$, for a photon traveling from the source to the observer. The number of orbits is then given by
@@
Based on the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.3 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.4. The function n(b)diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.3. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.4. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
@@ III.2 Transfer functions
We now study the emission from a thin disk viewed in a face-on orientation. In this case, we suppose that the thin disk is located on the Schwarzschild-Hernquist black hole’s equatorial plane and that its emission is isotropic in the rest frame of the emitting material. The specific intensity measured by a stationary observer at infinity, neglecting absorption and scattering, is given by
@@ III.3 Observational features
Building upon the framework established above, we can now compute and analyze the observed intensity profiles. The thin disk is the primary source of the observed flux for a distant observer, which is dependent only on the radial coordinate $r$. The face-on orientation provides the most general case Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). We thus consider three toy models for the emissivity profile of the thin disk Wang et al. (2022); Yang et al. (2023); Li and He (2021a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025):
[…]
@@
[…]
@@
[…]
@@
The parameters of the Hernquist DM halo have a direct and significant impact on the observational features, as can be seen by a comparison of Figs.6 and 7. The photon and lensing ring radii both enlarge with increasing parameter, while their contribution to the total flux diminishes. Furthermore, for Model 3, the diameter of the bright ring in the image appears to increase slightly with increasing DM parameters. As the Hernquist DM halo parameter grows, the maximum measured intensity increases for all three emission scenarios ($2.12\%$ for Model 1, $6.22\%$ for Model 2 and $5.31\%$ for Model 3). The peak intensity in Model 2 and 3 is significantly lower than in the corresponding Schwarzschild case Gralla et al. (2019). These findings provide theoretical predictions that could help interpret EHT observations and potentially constrain the presence and properties of a Hernquist DM halo around supermassive black holes.
@@ IV Shadows with spherical accretions
In addition to thin disk accretion, static and infalling spherical accretion flows can also form around Schwarzschild-Hernquist black holes. In this section, we further explore the influence of a Hernquist DM halo on the spacetime structure and the resulting observable properties under these spherical accretion models.
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadow properties of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. Employing ray-tracing techniques and numerical simulations, we have quantitatively analyzed three thin-disk accretion models (26)-(32), and two spherical accretion models Eqs.(37) and (43), revealing that the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ have a significant impact on the photon sphere radius and the observed shadow properties. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.6-11. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In extreme circumstances, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 9 and 11, the peak observed intensity is significantly diminished, giving the overall image of being substantially dimmer. The gravitational influence of the Hernquist DM halo is responsible for these alterations, modifying the effective potential around the black hole and the photon trajectory. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
@@
Importantly, our findings offer significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size—specifically, the positive correlation between the photon ring diameter and the Hernquist DM parameters—provides an observable diagnostic for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides an unambiguous prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, a future measurement of an anomalously large shadow could indicate the presence of a Hernquist DM halo, and our model could be used to interpret such a result. Conversely, the absence of significant shadow enlargement would place stringent constraints on the parameter space of viable Hernquist DM models.

### v2

@@ Abstract
The shadow of a black hole is critically dependent on the surrounding accreting matter. We investigate the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. Under spherical accretion, the Hernquist DM halo significantly enlarges the photon sphere. This results in an observable shadow that is approximately $2\%$ to $30\%$ larger than in the vacuum case, though this increase in size is accompanied by a considerable decrease in overall image brightness. Furthermore, the Doppler de-boosting effect in the infalling scenario produces a markedly darker image than its static counterpart. Our results demonstrate that the size and brightness profile of a black hole shadow provide a powerful observational tool to probe and constrain the distribution of dark matter in galactic centers.
@@ I Introduction
The observation of black hole shadows has entered a new era, establishing them as a ground-breaking tool for researching powerful gravitational forces. The fundamental nature of gravity has been further illuminated by analysis of the critical behaviour of photons close to black holes. Recent advances in the Event Horizon Telescope’s (EHT) imaging of the supermassive black holes $\text{M87}^{*}$ Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023) and $\text{SgrA}^{*}$ Ghez et al. (2005); Boehle et al. (2016); Akiyama et al. (2022a, b, c, d, e, f, 2024a, 2024b) have shown shadow features that not only support general relativity (GR) Hawking (1972); Bardeen et al. (1973a); Wald (1994, 2010) predictions but also provide a novel window of probing black holes. The black hole shadow, also known as the photon capture zone, is the centre dark area shown in these pictures Perlick and Tsupko (2022). This is encircles a compact, asymmetrical and dazzling circular structure. This phenomenon, known as the photon ring, results from light beams from infinity being deflected by the strong gravitational field as they approach the black hole. In backward ray-tracing methods, light rays that approach the critical impact parameter asymptotically converge to the bound photon orbit. The so-called photon sphere, which projects the shadow of the black hole, is made up of this confined photon orbit. An angular radius equation for the shadow of Schwarzschild black holes has been established by Synge and Luminet using a crucial impact parameter of $b_{p}=3\sqrt{3}M$ Synge (1966); Luminet (1979). The first image of a black hole surrounded by a thin accretion disk was analytically produced by the authors in Ref.Luminet (1979), displaying primary and secondary images that emerge beyond the black hole’s shadow. Bardeen later studied the D-shaped shadow of Kerr black holes, showing that the shape of the shadow can be changed by the spin of the black hole Bardeen et al. (1973b); Chandrasekhar (1998). Additionally, a great deal of research has been done on black hole shadows in higher-dimensional spacetime and other modified gravity theories Amarilla et al. (2010); Amarilla and Eiroa (2012); Amir et al. (2018); Mizuno et al. (2018); Eiroa and Sendra (2018); Banerjee et al. (2020); Bambi et al. (2019); Vagnozzi et al. (2023); Khodadi et al. (2024). In contrast to single shadows, the research has also suggested multiple shadows for wormholes and black holes Huang et al. (2016); Guerrero et al. (2021). While the finding of the EHT mainly relates to black holes in GR, it provides a great deal of opportunity to investigate other compact objects in theories other than GR.
@@
To investigate the main characteristics of black hole pictures and find possible indications of novel physics, it is helpful to set up a basic accretion model. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accret ion settings the shadow has resilient properties where its shape and size are determined by spacetime instead of accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Further investigations of optically thin and geometrically thin accretion models show that the size of the shadow is strongly related to the position of the accretion Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are always present in the bright areas outside the shadow. These characteristics make it possible to distinguish GR black holes from other compact objects or black holes in modified gravity theories through observation Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are extremely important for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@ II Light deflection and photon orbit
In this work, we model a Schwarzschild black hole immersed in a DM halo that is described by the Hernquist density profile Dehnen (1993); Mo et al. (2010),
@@
The radius $r_{p}$ and the impact parameter $b_{p}$ of the photon sphere are represented in Eqs.(15)-(16), respectively. Tab.1 displays numerical results for the event horizon radius $r_{h}$, photon sphere radius $r_{p}$, and impact parameter $b_{p}$ by choosing various values for $\rho_{c}$ and $r_{s}$. It is evident that as $\rho_{c}$ and $r_{s}$ increase, the event horizon radius $r_{h}$, the photon sphere radius $r_{p}$, and the impact parameter $b_{p}$ all rise, exceeding their values in the Schwarzschild case. This increase is a direct physical consequence of the additional mass-energy contributed by the DM halo, which enhances the spacetime curvature outside the event horizon. To maintain an unstable circular orbit within this stronger gravitational field, a photon must possess a larger angular momentum. For a distant observer, this corresponds to a larger critical impact parameter $b_{p}$, and the radius of the orbit itself $r_{p}$ is consequently shifted outwards. In essence, the DM halo enlarges the ability of the black hole for capturing photons.
@@ III.1 Direct emission, lensing ring and photon ring
According to Ref.Gralla et al. (2019), radiation with different impact parameters $b$ would cause the black hole to look differently to a distant observer when tracing rays backwards from an observer towards the black hole. We may use ray-tracing procedure to determine the optical appearance of a Schwarzschild-Hernquist black hole, by calculating the total deflection in azimuthal angle $\varphi$, for a photon traveling from the source to the observer. The number of orbits is then given by Gralla et al. (2019); Gralla and Lupsasca (2020)
@@
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.3 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.4. The function n(b)diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.3. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.4. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
@@ III.2 Transfer functions
We now study the emission from a thin disk viewed in a face-on orientation. In this case, we suppose that the thin disk is located on the Schwarzschild-Hernquist black hole’s equatorial plane and that its emission is isotropic in the rest frame of the emitting material. The specific intensity measured by a stationary observer at infinity, neglecting absorption and scattering, is given by Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a)
@@ III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the shadow and photon ring diameters with DM parameters, to be robust across different emission profiles.
Model 1 is designed to represent a standard, geometrically thin, optically thick accretion disk, analogous to the classical Novikov-Thorne model. The emissivity is sharply peaked at the innermost stable circular orbit (ISCO), which is the radius where gravitational energy is most efficiently released, and follows a rapid power-law decay outwards. This profile is characteristic of radiatively efficient accretion scenarios, often seen in luminous active galactic nuclei (AGN).
@@
Model 2 serves as an exploratory case to probe the observational appearance when the emission peak is located extremely close to the unstable photon orbits at $r_{p}$. While not a standard astrophysical scenario, this profile could hypothetically arise from localized energy release, such as magnetic reconnection events occurring near the photon sphere. Its primary purpose is to test the most extreme gravitational lensing effects on an emission source situated at the boundary of photon capture.
@@
Model 3 is chosen to represent a physically distinct scenario: a geometrically thick, optically thin, and radiatively inefficient accretion flow, such as an Advection-Dominated Accretion Flow (ADAF). This type of flow is believed to be present in low-luminosity AGN like $\text{M87}^{*}$ and $\text{SgrA}^{*}$. The emissivity profile is consequently broader, with a more gradual decay, reflecting emission that is spread over a much wider radial range rather than being concentrated at the inner edge.
@@
The parameters of the Hernquist DM halo have a direct and significant impact on the observational features, as can be seen by a comparison of Figs.6 and 7. The photon and lensing ring radii both enlarge with increasing parameter, while their contribution to the total flux diminishes. Furthermore, for Model 3, the diameter of the bright ring in the image appears to increase slightly with increasing DM parameters. As the Hernquist DM halo parameter grows, the maximum measured intensity increases for all three emission scenarios ($2.12\%$ for Model 1, $6.22\%$ for Model 2 and $5.31\%$ for Model 3).
@@ IV Shadows with spherical accretions
Having analyzed thin disk accretion, we now turn our attention to spherical accretion flows, which provide alternative astrophysical scenarios. We will investigate both static and infalling accretion models to determine how the Hernquist DM halo impacts the corresponding observational signatures.
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadows of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. Employing ray-tracing techniques and numerical simulations, we have quantitatively analyzed three thin-disk accretion models (26)-(32), and two spherical accretion models Eqs.(37) and (43), revealing that the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ have a significant impact on the photon sphere radius and the observed shadow properties. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.6-11. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In extreme circumstances, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 9 and 11, the peak observed intensity is significantly diminished, giving the overall image of being substantially dimmer. The gravitational influence of the Hernquist DM halo is responsible for these alterations, modifying the effective potential around the black hole and the photon trajectory. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
@@
Importantly, our researches offer significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, offers a powerful diagnostic for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides an unambiguous prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, a future measurement of an anomalously large shadow could indicate the presence of a Hernquist DM halo, and our model could be used to interpret such a result. Conversely, the absence of significant shadow enlargement would place stringent constraints on the parameter space of viable Hernquist DM models.